# Beyond Chain-of-Thought: A Survey of Chain-of-X Paradigms for LLMs

### v1

@@ Abstract
Chain-of-Thought (CoT) has been a widely adopted prompting method, eliciting impressive reasoning abilities of Large Language Models (LLMs). Inspired by the sequential thought structure of CoT, a number of Chain-of-X (CoX) methods have been developed to address various challenges across diverse domains and tasks involving LLMs. In this paper, we provide a comprehensive survey of Chain-of-X methods for LLMs in different contexts. Specifically, we categorize them by taxonomies of nodes, i.e., the X in CoX, and application tasks. We also discuss the findings and implications of existing CoX methods, as well as potential future directions. Our survey aims to serve as a detailed and up-to-date resource for researchers seeking to apply the idea of CoT to broader scenarios.
@@ 1 Introduction
Large Language Models (LLMs) have shown strong reasoning capabilities when prompted with the Chain-of-Thought (CoT) method Wei et al. (2022); Yao et al. (2024); Besta et al. (2024a). The essence of CoT is to decompose complex problems into sequences of intermediate subtasks Chu et al. (2023); Zhou et al. (2023). By handling these subtasks step by step, LLMs are able to focus on important details and assumptions, which substantially improves their performance across a wide range of reasoning tasks Huang and Chang (2023); Chu et al. (2023). Additionally, CoT’s intermediate steps offer a more transparent reasoning process, facilitating easier interpretation and evaluation of LLMs Yu et al. (2023b).
With the success of CoT, a number of Chain-of-X (CoX) methods have subsequently been developed Yu et al. (2023a). Extending beyond reasoning thoughts, recent CoX methods have constructed the chain with various components, such as Chain-of-Feedback Lei et al. (2023); Dhuliawala et al. (2023), Chain-of-Instructions Zhang et al. (2023d); Hayati et al. (2024), Chain-of-Histories Luo et al. (2024); Xia et al. (2024d), etc. These methods have been applied to tackle challenges in diverse tasks involving LLMs beyond reasoning, including multi-modal interaction Xi et al. (2023a); Zhang et al. (2024a), hallucination reduction Lei et al. (2023); Dhuliawala et al. (2023), planning with LLM-based agents Zhan and Zhang (2023); Zhang et al. (2024c), etc.
@@
Despite their growing prevalence, these CoX methods have not yet been collectively examined or categorized, leaving a gap in our understanding of their potential and nuances. To this end, this survey aims to offer a structured overview that captures the essence and diversity of CoX methods for further exploration and innovation.
@@
While several surveys have explored CoT Chu et al. (2023); Yu et al. (2023b); Besta et al. (2024b), they focus primarily on the reasoning thoughts of different structures, e.g., Chain-of-Thought as illustrated in Figure 1(a). In contrast, this paper focuses on the multifaceted component designs of Chain-of-X beyond reasoning thoughts as shown in Figure 1, offering insights of the CoT concept in broader domains. We present a comprehensive review by taxonomies of the X in CoX and tasks to which these methods are applied.
@@ 2 What is Chain-of-X?
CoT prompting is a methodology that substantially enhances the reasoning capabilities of LLMs. Introduced by Wei et al. (2022), CoT involves prompting LLMs with a structured format of <input, thoughts, output>, where ‘thoughts’ encompass coherent and intermediate natural language reasoning steps leading to the final answer. CoT’s effectiveness is most pronounced in tasks that require complex reasoning. Traditional few-shot learning methods often falter in such scenarios, as they tend to provide direct answers without the necessary intermediate steps. Rae et al. (2021) highlighted this limitation, noting the inadequacy of these methods with increased model size. In contrast, CoT prompting excels by incorporating intermediate reasoning steps. These steps guide the model through a logical progression, enhancing its capability to tackle complex problems, such as those involving arithmetic, commonsense, and symbolic reasoning Wang et al. (2023d); Lyu et al. (2023). The essence of CoT lies in its strategy to tackle complex problems by breaking them down into manageable intermediate steps Zhou et al. (2023). Kojima et al. (2022) have also demonstrated strong performance of zero-shot CoT by prompting “Let’s think step by step.”. The explicit reasoning steps also provide a transparent pathway for the model’s thought process, allowing for further evaluations and corrections Yu et al. (2023b).
@@
Inspired by the nature of the sequential breakdown of CoT, a substantial number of CoX methods have been developed recently Yu et al. (2023a). Here, we define CoX as a generalization of the CoT method for diverse tasks beyond LLM reasoning. We refer to the X in CoX as the ‘node’ of the chain structure. Beyond the thoughts in CoT prompts, the X in CoX can take various forms tailored to specific tasks, including intermediates (§3.1), augmentation (§3.2), feedback (§3.3), and even models (§3.4), as illustrated in Figure 1. We summarize the types of nodes in existing CoX methods in Figure 2. The idea of CoX is to construct a sequence of problem-related components that either compositionally contribute to the solution or iteratively refine the outputs for a complex task. Similarly, we define a structured format for CoX as <input, X1, , Xn, output> where is the length of the chain. Note that this format extends beyond prompting strategies like CoT and can be adapted to a variety of algorithmic frameworks or structures for diverse tasks involving LLMs. For instance, Chain-of-Verification Dhuliawala et al. (2023) is a hallucination reduction framework that employs an LLM to generate initial responses, composes a sequence of verification questions, and revises its previous responses based on these questions. In addition to hallucination reduction, CoX methods have been applied to a variety of tasks, as shown in Figure 2, including multi-modal interaction (§4.1), factuality & safety (§4.2), multi-step reasoning (§4.3), instruction following (§4.4), LLMs as Agents (§4.5), and evaluation tools (§4.6).
for tree=
font=,
draw=myblue, semithick, rounded corners,
minimum height = 1.5ex,
minimum width = 2em,
anchor = west,
grow = east,
forked edge, s sep = 0.8mm, l sep = 3.5mm, fork sep = 2mm,
[A Survey of Chain-of-X, rotate=90, anchor=center
[Taxonomy of Tasks (§4), fit=band, text width=1cm
[Evaluation Tools
(§4.6), text width=2.1cm, l sep = 2mm
[E.g., CoUtterances Bhardwaj and Poria (2023), CoImages Meng et al. (2023), CoFeedback Ahn and Shin (2024), text width=10.9cm, fill=lightblue]
[LLMs as Agents
(§4.5), text width=2.1cm, l sep = 2mm
[E.g., CoActiona Zhan and Zhang (2023), CoSummarization Ma et al. (2023), Co3DThought Yamada et al. (2024), CoActionThought Zhang et al. (2024c), CoAbstraction Gao et al. (2024), CoContacts Xiao et al. (2024a), text width=10.9cm, fill=lightblue]
[Instruction Following (§4.4), text width=2.1cm, l sep = 2mm
[E.g., CoTask Li et al. (2024b), CoInstructions Hayati et al. (2024), CoModality Zhang et al. (2023a), text width=10.9cm, fill=lightblue]
[Multi-Step Reasoning (§4.3), text width=2.1cm, l sep = 2mm
[E.g., CoT Wei et al. (2022), CoKnowledgeb Wang et al. (2023c), CoMemory Hu et al. (2023a), CoReference Kuppa et al. (2023), CoQuery Xu et al. (2023), CoLogic Servantez et al. (2024), CoFeedback Ahn and Shin (2024), text width=10.9cm, fill=lightblue]
[Factuality & Safety (§4.2), text width=2.1cm
[Alignment, text width=1.3cm, l sep = 2mm
[E.g., CoUtterances Bhardwaj and Poria (2023), CoHindsight Liu et al. (2024a), text width=9.0cm, fill=lightblue]
[Hallucination Reduction, text width=1.3cm, l sep = 2mm
[E.g., CoNLI Lei et al. (2023), CoVerification Dhuliawala et al. (2023), CoKnowledgea Li et al. (2024a), CoNote Yu et al. (2023a), CoQuestion Huang et al. (2024), CoActionb Pan et al. (2024), text width=9.0cm, fill=lightblue]
[Multi-Modal Interaction (§4.1), text width=2.1cm
[Text-Speech, text width=1.3cm, l sep = 2mm
[E.g., CoInformation Zhang et al. (2024a), CoModality Zhang et al. (2023a), text width=9.0cm, fill=lightblue]
[Text-Code, text width=1.3cm, l sep = 2mm
[E.g., CoRepair Wang et al. (2023b), CoCode Li et al. (2023a), CoSimulation La Malfa et al. (2024), text width=9.0cm, fill=lightblue]
[Text-Table, text width=1.3cm, l sep = 2mm
[E.g., CoCommand Zha et al. (2023), CoTable Wang et al. (2024), text width=9.0cm, fill=lightblue]
[Text-Image, text width=1.3cm, l sep = 2mm
[E.g., CoLook Xi et al. (2023a), CoSpot Liu et al. (2024b), CoReasoning Uehara et al. (2024), text width=9.0cm, fill=lightblue]
[Taxonomy of Nodes (§3), fit=band, text width=1cm
[Chain-of-Models
(§3.4), text width=2.1cm, l sep = 2mm
[E.g., CoExperts Xiao et al. (2024b), CoDiscussion Tao et al. (2024), CoLoRA Xia et al. (2024a), text width=10.9cm, fill=lightblue]
[Chain-of-Feedback (§3.3), text width=2.1cm
[Self-Refine, text width=1.3cm, l sep = 2mm
[E.g., CoVerification Dhuliawala et al. (2023), CoSelfRevisions Le et al. (2024), text width=9.0cm, fill=lightblue]
[External
Feedback, text width=1.3cm, l sep = 2mm
[E.g., CoRepair Wang et al. (2023b), CoHindsight Liu et al. (2024a), text width=9.0cm, fill=lightblue]
[Chain-of-Augmentation (§3.2), text width=2.1cm
[Others, text width=1.3cm, l sep = 2mm
[E.g., CoReference Kuppa et al. (2023), CoDictionary Lu et al. (2023), CoMemory Hu et al. (2023a), text width=9.0cm, fill=lightblue]
[Retrievals, text width=1.3cm, l sep = 2mm
[E.g., CoQuery Xu et al. (2023), CoKnowledgea Li et al. (2024a), CoQuestion Huang et al. (2024), text width=9.0cm, fill=lightblue]
[Histories, text width=1.3cm, l sep = 2mm
[E.g., CoOpinion Do et al. (2023), CoHistorya Luo et al. (2024), CoHistoryb Xia et al. (2024d), text width=9.0cm, fill=lightblue]
[Instructions, text width=1.3cm, l sep = 2mm
[E.g., CoInstructEditing Zhang et al. (2023d), CoInstructions Hayati et al. (2024), text width=9.0cm, fill=lightblue]
[Chain-of-Intermediates (§3.1), text width=2.1cm
[Knowledge
Composition, text width=1.3cm, l sep = 2mm
[E.g., CoSymbol Hu et al. (2023b), CoKnowledgeb Wang et al. (2023c), CoManipulation Qi et al. (2024), CoSimulation La Malfa et al. (2024), CoSpot Liu et al. (2024b), text width=9.0cm, fill=lightblue]
[Problem
Decomposition, text width=1.3cm, l sep = 2mm
[E.g., CoT Wei et al. (2022), CoCode Li et al. (2023a), CoTable Wang et al. (2024), CoLogic Servantez et al. (2024), CoEvent Bao et al. (2024), CoInteraction Han et al. (2024), text width=9.0cm, fill=lightblue]
@@ 3.1 Chain-of-Intermediates
Building on the concept of utilizing explicit intermediate steps, a natural evolution of CoT involves replacing reasoning thoughts with other types of intermediate components, i.e., Chain-of-Intermediates. Based on the primary focuses, we further divide them into the following subtypes.
@@
In problem decomposition, the intermediate steps consist of manageable subtasks derived from an original complex problem. This approach is exemplified by the classic Chain-of-Thought prompting Wei et al. (2022). Extending this further, Li et al. (2023a) introduce Chain-of-Code, which segments a task into programmatic subtasks, enhancing the reasoning process through simulated code outputs. Similarly, Wang et al. (2024) have developed the Chain-of-Table framework. This framework restructures complex tables into question-specific formats via a sequence of strategic operations, making the data more accessible and tailored to the inquiry. Moreover, the Chain-of-Logic, introduced by Servantez et al. (2024), applies a logical decomposition to rule-based reasoning tasks, transforming them into a series of logical expressions. The methodological breakdown facilitates clearer reasoning pathways. These decomposition methods are also echoed in Chain-of-Event Han et al. (2024), which simplify multi-document summarization into discrete and manageable event extraction tasks, significantly enhancing quality and reducing potential errors.
@@
In knowledge composition, the primary goal of the intermediate steps is not on simplification but on the accumulation of relevant information and evidence. This approach aims to enrich the solution with a depth of understanding and details. For instance, Hu et al. (2023b) propose Chain-of-Symbol method that meticulously collects spatial relations during spatial planning tasks, enhancing the model’s precision and effectiveness. Likewise, La Malfa et al. (2024) adopt Chain-of-Simulation prompting to ensures that each step in code execution is informed by program traces, thereby avoiding memorization pitfalls. Wang et al. (2023c) take a similar approach with Chain-of-Knowledge, extracting crucial pieces of evidence at each step to support more grounded and reliable question-answering sessions. This technique is particularly effective in fostering a deeper understanding of the queried material. In visual tasks, methods like Chain-of-Spot Liu et al. (2024b) and Chain-of-Reasoning Uehara et al. (2024) assist vision-language models in focusing on specific image details, which is crucial for tasks requiring detailed visual evidences. Through these evidence-rich methods, LLMs achieve a comprehensive and nuanced understanding of complex scenarios, leading to higher quality outputs.

### v3

@@ Abstract
Chain-of-Thought (CoT) has been a widely adopted prompting method, eliciting impressive reasoning abilities of Large Language Models (LLMs). Inspired by the sequential thought structure of CoT, a number of Chain-of-X (CoX) methods have been developed to address challenges across diverse domains and tasks. In this paper, we provide a comprehensive survey of Chain-of-X methods for LLMs in different contexts. Specifically, we categorize them by taxonomies of nodes, i.e., the X in CoX, and application tasks. We also discuss the findings and implications of existing CoX methods, as well as potential future directions. Our survey aims to serve as a detailed and up-to-date resource for researchers seeking to apply the idea of CoT to broader scenarios.
@@ 1 Introduction
Large Language Models (LLM) have shown strong reasoning capabilities when prompted with the Chain-of-Thought (CoT) method Wei et al. (2022). The essence of CoT is to decompose complex problems into sequences of intermediate subtasks Chu et al. (2024). By handling these subtasks step by step, LLMs are able to focus on important details and assumptions, which substantially improves the performance across a wide range of reasoning tasks Huang and Chang (2023); Chu et al. (2024). CoT’s intermediate steps can also offer a more transparent reasoning process, facilitating easier interpretation and evaluation of LLMs Yu et al. (2023b).
With the success of CoT, a number of Chain-of-X (CoX) methods have subsequently been developed Yu et al. (2023a). Extending beyond reasoning thoughts, recent CoX methods have constructed the chain with various components, such as Chain-of-Feedback Lei et al. (2023); Dhuliawala et al. (2023), Chain-of-Instructions Zhang et al. (2023d); Hayati et al. (2024), Chain-of-Histories Luo et al. (2024); Xia et al. (2024e), etc. These methods have been applied to diverse tasks involving LLMs beyond reasoning, including multi-modal interaction Xi et al. (2023a); Zhang et al. (2024a), hallucination reduction Lei et al. (2023); Dhuliawala et al. (2023), planning with LLM-based agents Zhang and Zhang (2024); Zhang et al. (2024b), etc.
@@
Despite their growing prevalence, these CoX methods have not yet been collectively examined or categorized, leaving a gap in our understanding of their potential. This survey offers a structured overview capturing CoX’s essence and diversity for further exploration and innovation.
@@
While several surveys have explored CoT Chu et al. (2024); Yu et al. (2023b); Besta et al. (2024), they focus primarily on the reasoning thoughts of different structures, e.g., Chain-of-Thought as illustrated in Figure 1(a). In contrast, this paper focuses on the multifaceted component designs of Chain-of-X beyond reasoning thoughts as shown in Figure 1, offering insights of the CoT concept in broader domains. We present a comprehensive review by taxonomies of the X in CoX and tasks to which these methods are applied.
@@ 2 What is Chain-of-X?
CoT prompting is a methodology that substantially enhances the reasoning capabilities of LLMs. Introduced by Wei et al. (2022), CoT involves prompting LLMs with a structured format of <input, thoughts, output>, where ‘thoughts’ encompass coherent and intermediate natural language reasoning steps leading to the final answer. CoT’s effectiveness is most pronounced in tasks that require complex reasoning. Traditional few-shot learning methods often falter in such scenarios, as they tend to provide direct answers without rationales. In contrast, CoT prompting excels by breaking down a complex task into manageable intermediate steps. These steps guide the model through a logical progression, enhancing its capability to tackle complex problems such as arithmetic, commonsense, and symbolic reasoning Wang et al. (2023f); Lyu et al. (2023). Additionally, Kojima et al. (2022) have also demonstrated strong performance of zero-shot CoT by prompting “Let’s think step by step.”. The explicit reasoning steps also provide a transparent pathway for the model’s thought process, allowing for further evaluations and corrections Yu et al. (2023b).
@@
Inspired by the nature of the sequential breakdown, a substantial number of CoX methods have been developed recently Yu et al. (2023a). Here, we define CoX as a generalization of the CoT method for diverse tasks beyond LLM reasoning. We refer to the X in CoX as the ‘node’ of the chain structure. Beyond the thoughts in CoT prompts, the X in CoX can take various forms tailored to specific tasks, including intermediates (§3.1), augmentation (§3.2), feedback (§3.3), and even models (§3.4), as illustrated in Figure 1. We summarize the types of nodes in existing CoX methods in Figure 2. The idea of CoX is to construct a sequence of problem-related components that either compositionally contribute to the solution or iteratively refine the outputs. We define a similar structured format for CoX as <input, X1, , Xn, output> where is the length of the chain. Note that this format extends beyond prompting-based strategies like CoT and can be adapted to a variety of algorithmic frameworks or structures for diverse tasks involving LLMs. For instance, Chain-of-Verification Dhuliawala et al. (2023) is a hallucination reduction framework that employs an LLM to generate initial responses, composes a sequence of verification questions, and revises its previous responses based on these questions. In addition to hallucination reduction, CoX methods have been applied to a variety of tasks, as shown in Figure 2, including multi-modal interaction (§4.1), factuality & safety (§4.2), multi-step reasoning (§4.3), instruction following (§4.4), LLMs as Agents (§4.5), and evaluation tools (§4.6).
for tree= font=, draw=myblue, semithick, rounded corners, minimum height = 1.ex, minimum width = 2em, anchor = west, grow = east, forked edge, s sep = 0.5mm, l sep = 2mm, fork sep = 1mm, [A Survey of Chain-of-X, rotate=90, anchor=center [Taxonomy of Tasks (§4), fit=band, text width=0.9cm [Evaluation Tools (§4.6), text width=2.61cm, l sep = 2mm [E.g., CoUtterances Bhardwaj and Poria (2023), BadChain Xiang et al. (2024), CoFeedback Ahn and Shin (2024), text width=10.6cm, fill=lightblue] ] [LLMs as Agents (§4.5), text width=2.61cm, l sep = 2mm [E.g., CoActiona Zhang and Zhang (2024), CoActionThought Zhang et al. (2024b), CoContacts Xiao et al. (2024a), text width=10.6cm, fill=lightblue] ] [Instruction Following (§4.4), text width=2.61cm, l sep = 2mm [E.g., CoTask Li et al. (2024b), LogiCoT Liu et al. (2023), CoImagination Zhou et al. (2024), text width=10.6cm, fill=lightblue] ] [Multi-Step Reasoning (§4.3), text width=2.61cm, l sep = 2mm [E.g., CoDensity Adams et al. (2023), CoLogic Servantez et al. (2024), CoEvent Bao et al. (2024), text width=10.6cm, fill=lightblue] ] [Factuality & Safety (§4.2), text width=2.61cm [Alignment, text width=1.35cm, l sep = 1mm [E.g., CoUtterances Bhardwaj and Poria (2023), CoHindsight Liu et al. (2024a), text width=8.9cm, fill=lightblue] ] [Hallucination, text width=1.35cm, l sep = 1mm [E.g., CoVerification Dhuliawala et al. (2023), CoKnowledgea Li et al. (2024a), text width=8.9cm, fill=lightblue] ] ] [Multimodal Interaction (§4.1), text width=2.61cm [Text-Speech, text width=1.35cm, l sep = 1mm [E.g., CoInformation Zhang et al. (2024a), CoModality Zhang et al. (2023a), text width=8.9cm, fill=lightblue] ] [Text-Code, text width=1.35cm, l sep = 1mm [E.g., CoRepair Wang et al. (2023b), CoCode Li et al. (2023a), CoSimulation La Malfa et al. (2024), text width=8.9cm, fill=lightblue] ] [Text-Table, text width=1.35cm, l sep = 1mm [E.g., CoCommand Zha et al. (2023), CoTable Wang et al. (2024b), text width=8.9cm, fill=lightblue] ] [Text-Image, text width=1.35cm, l sep = 1mm [E.g., MMCoT Zhang et al. (2023e), CoLook Xi et al. (2023a), CoSpot Liu et al. (2024c), text width=8.9cm, fill=lightblue] ] ] ] [Taxonomy of Nodes (§3), fit=band, text width=0.9cm [Chain-of-Models (§3.4), text width=2.61cm, l sep = 2mm [E.g., CoExperts Xiao et al. (2024b), ChatEval Chan et al. (2024), CoDiscussion Tao et al. (2024), text width=10.6cm, fill=lightblue] ] [Chain-of-Feedback (§3.3), text width=2.61cm [Self Feedback, text width=1.35cm, l sep = 1mm [E.g., Self-Refine Madaan et al. (2023), CoVerification Dhuliawala et al. (2023), text width=8.9cm, fill=lightblue] ] [Ext. Feedback, text width=1.35cm, l sep = 1mm [E.g., Chain-of-3DThought Yamada et al. (2024), CoHindsight Liu et al. (2024a), text width=8.9cm, fill=lightblue] ] ] [Chain-of-Augmentation (§3.2), text width=2.61cm [Others, text width=1.35cm, l sep = 1mm [E.g., CoReference Kuppa et al. (2023), CoDictionary Lu et al. (2023), CoMemory Hu et al. (2023a), text width=8.9cm, fill=lightblue] ] [Tools, text width=1.35cm, l sep = 1mm [E.g., ChatCoT Chen et al. (2023b), MultiToolCoT Inaba et al. (2023), CoAbstract Gao et al. (2024), text width=8.9cm, fill=lightblue] ] [Retrievals, text width=1.35cm, l sep = 1mm [E.g., ReAct Yao et al. (2023), CoQuery Xu et al. (2024), CoKnowledgea Li et al. (2024a), text width=8.9cm, fill=lightblue] ] [Histories, text width=1.35cm, l sep = 1mm [E.g., CoOpinion Do et al. (2023), CoHistorya Luo et al. (2024), CoHistoryb Xia et al. (2024e), text width=8.9cm, fill=lightblue] ] [Instructions, text width=1.35cm, l sep = 1mm [E.g., CoInstructEditing Zhang et al. (2023d), CoInstructions Hayati et al. (2024), text width=8.9cm, fill=lightblue] ] ] [Chain-of-Intermediates (§3.1), text width=2.61cm [Know. Comp., text width=1.35cm, l sep = 1mm [E.g., CoKnowledgeb Wang et al. (2023d), CueCoT Wang et al. (2023c), CoSpot Liu et al. (2024c), text width=8.9cm, fill=lightblue] ] [Prob. Decomp., text width=1.35cm, l sep = 1mm [E.g., CoT Wei et al. (2022), Least-to-Most Zhou et al. (2023), CoCode Li et al. (2023a), text width=8.9cm, fill=lightblue] ] ] ] ]
@@ 3.1 Chain-of-Intermediates
Building on the concept of utilizing intermediate steps, a natural evolution of CoT involves generalizing reasoning thoughts to other types of intermediate components. Based on the primary focuses, we further divide them into the following subtypes.
@@
In problem decomposition, the intermediate steps consist of manageable subtasks derived from an original complex problem, which is exemplified by the classic Chain-of-Thought prompting Wei et al. (2022) for reasoning tasks. To overcome the challenge of easy-to-hard generalization, Zhou et al. (2023) further introduce Least-to-Most prompting which breaks down a complex problem into simpler subtasks and solves them in sequence. Extending beyond natural languages, Chain-of-Code Li et al. (2023a) takes advantage of the syntactic structure and precise computation of code by segmenting a complex task into programmatic subtasks, enhancing the reasoning process through simulated code outputs. Similarly motivated by computation accuracy, Chain-of-Logic Servantez et al. (2024) applies a logical decomposition transforming rule-based reasoning tasks into a series of simple logical expressions.
While such decomposition is widely applied in reasoning tasks, the concept is also echoed in other tasks. For example, Chain-of-Event Han et al. (2024) simplifies multi-document summarization into discrete event extraction tasks, significantly enhancing the summarization quality and factuality. Chain-of-Table Wang et al. (2024b) restructures complex tables into question-specific formats via a sequence of strategic operations, making the data more accessible and tailored to the inquiry.
@@
In knowledge composition, the primary goal of the intermediate steps is not simplification but the accumulation of relevant information. This approach aims to enrich the solution with a depth of understanding and details. For instance, to handle unfactual rationales generated with CoT prompting, Wang et al. (2023d) propose Chain-of-Knowledgeb to elicit LLMs to generate explicit knowledge evidence at each reasoning step for more grounded question-answering. Similarly in dialogue systems, CueCoT Wang et al. (2023c) collects linguistic cues with intermediate steps to capture contextual user status for more personalized and engaging conversation.
Besides natural language tasks, this technique is also useful in knowledge-intensive visual tasks that require capturing specific visual details. For example, Chain-of-Spot Liu et al. (2024c) and Chain-of-Reasoning Uehara et al. (2024) enable vision-language models to focus on key regions of interest, improving reasoning performance with detailed visual evidences. Likewise, CCoT Mitra et al. (2024) utilizes scene-graph representations to extract compositional knowledge from a large multimodal model, which is further used to facilitate its own response generation on vision-language tasks.